OpenStudy (anonymous):

On a street with 24 houses, 9 houses have fewer than 6 rooms, 8 houses have more than 7 rooms, and 4 house have more than 8 room. What is the total number of houses on the street that are either 6-, 7- or 8-room houses? 8 9 11 13 14

OpenStudy (e.mccormick):

So, how far did you get with this?

OpenStudy (anonymous):

i had got 11 and it wqas wrong

OpenStudy (debbieg):

I was in the process of replying on your OTHER post of this same question when you closed it. I'll try again.

OpenStudy (e.mccormick):

They key here is you are lookng for things that are from 6 to 8, right? So what in the information excludes parts of that?

OpenStudy (debbieg):

Starting with 24 houses, 9 have FEWER than 6 rooms. So those are NOT in the category of 6, 7 or 8 room houses, so eliminate those. That leaves 24 - 9 =15 houses to consider.

OpenStudy (debbieg):

Since 8 houses have MORE THAN 7 (includes the 8-room houses that we want to count), and 4 houses have MORE THAN 8 (which we DONT want to include in our count), that tells us to eliminate how many more, from the 15 we are now down to?

OpenStudy (anonymous):

14

OpenStudy (debbieg):

How did you get 14?

OpenStudy (debbieg):

If 8 have more than 7, and 4 have more than 8, then how many of the 8 that have more than 7, do we need to eliminate from our count?

OpenStudy (debbieg):

We WANT the 8-room houses, but not the more-than-8-room houses. If 8 houses have more than 7, and 4 have more than 8, then how many MORE houses should we EXCLUDE from our count?

OpenStudy (debbieg):

hmmmm.... you say that you answered 11 and it was wrong?

OpenStudy (anonymous):

yea

OpenStudy (debbieg):

Well, lets see what @e.mccormick says, because I think that's the correct answer too. lol

OpenStudy (e.mccormick):



OpenStudy (e.mccormick):

Becaue you are looking for 6, 7, and 8, you need to see what eliminates more than 8 and less than 6. Any data that talks about more than 7 is extraneous.

OpenStudy (e.mccormick):

Oh, and there are exactly 4 houses with 8 rooms, but that doe not matter.

OpenStudy (anonymous):

it's 11

OpenStudy (debbieg):

You start with 24. You eliminate: 9 that have fewer than 6 rooms 4 that have more than 8 rooms So that's 9+4=13 24 - 13 = the ones that have 6 - 8 rooms.

OpenStudy (e.mccormick):

Yes. There are 11. So if you answered that, it is correct.

OpenStudy (e.mccormick):

So yah, you are not crazy and 24-9-4=11 is still true.